Find the length of the diagonal of a rectangle whose length and breadth are 4m and 3m respectively.

Answer by checkley79(3341)

YOU HAVE A RIGHT TRIANGLE WITH SIDES OF 3 & 4 M.
A^2+B^2=C^2
3^+4^=C^2
9+16=C^2
25=C^2
C=SQRT25
C=5 M ANS. FOR THE DIAGONAL.